Payment of 700.00 every month for one (1) year that will start at the end of the third month. What is the period of deferral?


1
Expert's answer
2022-01-19T08:22:05-0500

Solution:

Given:

"t=\frac{3}{12}"

"n=12"

"k=nt-1"

"k=(\frac{3}{12}×12)-1=2"